# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v3

@@ A.5 Detailed results
Table 8: Full comparison of $\ell_{1}$-ERT. $\ell_{1}$-ERT ($\downarrow$) on all benchmarks with all baselines.
@@
Table 9: Full comparison of $\ell_{2}$-ERT. $\ell_{2}$-ERT ($\downarrow$) on all benchmarks with all baselines.
@@ Appendix B Proofs
denote the corresponding expected pinball loss. Let $q^{*}=q_{\tau}(x)$ be the true $\tau$-quantile, i.e., $F(q^{*})=\tau$, and let $\hat{q}=\hat{q}_{\tau}(x)$.
@@
Since $q^{*}$ is a minimizer of $\mathcal{L}$, we have $F(q^{*})-\tau=0$.
@@
$\mathcal{L}(\hat{q})-\mathcal{L}(q^{*})=\int_{q^{*}}^{\hat{q}}\bigl(F(z)-\tau\bigr)dz.$ (40)
@@
Without loss of generality, assume $\hat{q}\geq q^{*}$ (the case $\hat{q}<q^{*}$ follows by symmetry). Define $g(z)=F(z)-\tau$, so that $g(q^{*})=0$, $g(z)\geq 0$ for $z\geq q^{*}$, and $g^{\prime}(z)=f(z)\leq L_{F}$ almost everywhere.
@@
$H(q)=\int_{q^{*}}^{q}g(z)dz-\frac{1}{2L_{F}}g(q)^{2},\qquad q\geq q^{*}.$
For almost every $q\geq q^{*}$, we have
@@
where the inequality follows from $g(q)\geq 0$ and $f(q)\leq L_{F}$. Since $H(q^{*})=0$ and $H$ is non-decreasing on $q^{*},\infty)$, it follows that $H(\hat{q})\geq 0$, i.e.,
$\int_{q^{*}}^{\hat{q}}\bigl(F(z)-\tau\bigr)dz\geq\frac{1}{2L_{F}}\bigl(F(\hat{q})-\tau\bigr)^{2}.$ (41)
@@
$\mathcal{L}(\hat{q})-\mathcal{L}(q^{*})\geq\frac{1}{2L_{F}}\bigl(F(\hat{q})-\tau\bigr)^{2},$
@@
$\bigl|F(\hat{q})-\tau\bigr|\leq\sqrt{2L_{F}\bigl(\mathcal{L}(\hat{q})-\mathcal{L}(q^{*})\bigr)}.$
@@
$=f_{S\mid X}(q_{\tau}(x))^{2}\epsilon_{q}(x)^{2}+\frac{1}{6}G^{\prime\prime\prime}(\xi_{S,1})\epsilon_{q}(x)^{3},$
@@ B.3.1 Preliminary: fast rates via local Rademacher complexity
Without additional curvature assumptions, global complexity bounds typically yield the slow rate $\mathcal{E}_{q}(n)=O_{\mathbb{P}}(n^{-1/4})$. To obtain fast rates, we invoke the theory of local Rademacher complexity [Bartlett et al., 2005].
@@ B.3.4 Bounding the weight estimation error (Term II)
Next, we leverage Assumption 5.1.3 ($1/q^{\prime}_{\tau}(x)\leq B_{w}$) and to make the discussion meaningful, we must assume $\delta$ is small enough ($\delta\leq\sqrt{12/B_{w}B_{q}^{\prime\prime\prime}}$), so that:
@@
We note that we should consider two $g^{*}$s, i.e., $q_{\beta}$ for $\beta\in\{\tau\pm\delta\}$. Moreover, for notational convenience, we will denote $C_{\mathrm{norm}}=C_{\mathrm{H}}R^{\frac{d}{2s+d}}$, and $\nu=2s/(2s+d)$.
@@
For classic statistical regression models, e.g., parametric model or Reproducing Kernel Hilbert Space (RKHS) regression with smooth kernel (e.g., Gaussian kernel), we have $\nu\approx 1$. Even for the complex ReLU multi-layer perceptrons with bounded weights and depth, functions in $\mathcal{G}$ are piecewise linear and hence belong to the Hölder class $C^{1}(\mathcal{X})$ with norm bounded by a constant depending on the network architecture. In this case, Assumption B.3 applies with $s=1$, yielding the exponent $2/(2+d)$ in (67). For other smoother activation functions such as $\tanh$, $s$ would be higher.
@@
$\hat{\Delta}_{\delta}(x)$ $\geq\Delta_{\delta}(x)-|\Delta_{\delta}(x)-\hat{\Delta}(x)|$ (68)
$\geq\Delta_{\delta}(x)-(|\hat{q}_{\tau+\delta}(x)-q_{\tau+\delta}(x)|+|\hat{q}_{\tau-\delta}(x)-q_{\tau-\delta}(x)|)$
$\geq\frac{1}{2B_{w}}-(\|\hat{q}_{\tau+\delta}-q_{\tau+\delta}\|_{L_{\infty}}+\|\hat{q}_{\tau-\delta}-q_{\tau-\delta}\|_{L_{\infty}})$
@@
$C_{\text{norm}}\left(C_{\text{fast}}\mathfrak{R}_{n}(\mathcal{G})\right)^{\nu}\leq\frac{1}{2B_{w}},$ (69)
which translates to the a substantial lower bound on sample size $n$.
@@
Hence, we can applying Bernstein’s inequality [Bach, 2024] to transform the empirical norm into population norm, with probability at least $1-\zeta$:
@@ B.3.5 Optimal bandwidth and final bound
$c_{2}=4B_{w}^{2}\sqrt{2\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}}.$ (76)
@@
$U(\delta^{\star})/M_{\rho,\mathcal{G}}$ $=3\left(\frac{c_{1}}{2}\right)^{1/3}c_{2}^{2/3}$ (77)
$=C_{\mathrm{const}}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}\left(2\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}\right)^{1/3},$
@@
$+2M_{\rho,\mathcal{G}}C_{\mathrm{const}}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}\left(2\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}\right)^{1/3}$
@@ C.1 Gap between $F_{S\mid X}(\hat{q}(x))$ and $\pi(x)$
In this part, we supplement the results proved in Plassier et al. [2025a].
@@
Here, we adopt a slightly stronger  [^2] We assume the existence of PDF, while in  Plassier et al. [2025a], the continuity of CDF is assumed. version than Plassier et al. [2025a] and suppress complex notations to enhance the readability. Assumption C.1 essentially characterizes the quality of quantile estimator, i.e., to what extent the quantile estimator removes the covariate-dependency on scores.
Theorem C.2 (Theorem 2 in Plassier et al. [2025a]).
@@
We note that the conformalization step is still substantial when $n$ is small (e.g., smaller than 100), while the $\Lambda$ would also decrease with the increase of the quality of the quantile estimator (and originally, the size of samples used for training the quantile estimator). Hence, with a reasonably large sample size of $\mathcal{D}_{\text{cal}}$, we can focus on characterizing the finite-sample performance of our method through analyzing the MSE of $F_{S\mid X}\left(\hat{q}_{\tau}(x)\right)$, and naturally the surrogate, expected risk of density-weighted pinball loss. We now present our main assumptions and theorem as follows.

### v4

@@ A.5 Detailed Results
Table 8: Full comparison of $L_{1}$-ERT. $L_{1}$-ERT ($\downarrow$) on all benchmarks with all baselines.
@@
Table 9: Full comparison of $L_{2}$-ERT. $L_{2}$-ERT ($\downarrow$) on all benchmarks with all baselines.
@@ Appendix B Proofs
denote the corresponding expected pinball loss. Let $q^{\star}=q_{\tau}(x)$ be the true $\tau$-quantile, i.e., $F(q^{\star})=\tau$, and let $\hat{q}=\hat{q}_{\tau}(x)$.
@@
Since $q^{\star}$ is a minimizer of $\mathcal{L}$, we have $F(q^{\star})-\tau=0$.
@@
$\mathcal{L}(\hat{q})-\mathcal{L}(q^{\star})=\int_{q^{\star}}^{\hat{q}}\bigl(F(z)-\tau\bigr)dz.$ (40)
@@
Without loss of generality, assume $\hat{q}\geq q^{\star}$ (the case $\hat{q}<q^{\star}$ follows by symmetry). Define $g(z)=F(z)-\tau$, so that $g(q^{\star})=0$, $g(z)\geq 0$ for $z\geq q^{\star}$, and $g^{\prime}(z)=f(z)\leq L_{F}$ almost everywhere.
@@
$H(q)=\int_{q^{\star}}^{q}g(z)dz-\frac{1}{2L_{F}}g(q)^{2},\qquad q\geq q^{\star}.$
For almost every $q\geq q^{\star}$, we have
@@
where the inequality follows from $g(q)\geq 0$ and $f(q)\leq L_{F}$. Since $H(q^{\star})=0$ and $H$ is non-decreasing on $q^{\star},\infty)$, it follows that $H(\hat{q})\geq 0$, i.e.,
$\int_{q^{\star}}^{\hat{q}}\bigl(F(z)-\tau\bigr)dz\geq\frac{1}{2L_{F}}\bigl(F(\hat{q})-\tau\bigr)^{2}.$ (41)
@@
$\mathcal{L}(\hat{q})-\mathcal{L}(q^{\star})\geq\frac{1}{2L_{F}}\bigl(F(\hat{q})-\tau\bigr)^{2},$
@@
$\bigl|F(\hat{q})-\tau\bigr|\leq\sqrt{2L_{F}\bigl(\mathcal{L}(\hat{q})-\mathcal{L}(q^{\star})\bigr)}.$
@@
$=\frac{1}{2}f_{S\mid X}(q_{\tau}(x))^{2}\epsilon_{q}(x)^{2}+\frac{1}{6}G^{\prime\prime\prime}(\xi_{S,1})\epsilon_{q}(x)^{3},$
@@ B.3.1 Preliminary: fast rates via local Rademacher complexity
Without additional curvature assumptions, global complexity bounds typically yield the slow rate $\mathcal{E}_{q}(n)=O_{\mathbb{P}}(n^{-1/4})$. To obtain fast rates, we invoke the theory of local Rademacher complexity (Bartlett et al., 2005).
@@ B.3.4 Bounding the weight estimation error (Term II)
Next, we leverage Assumption 5.1.3 ($1/q^{\prime}_{\tau}(x)\leq B_{w}$) and to make the discussion meaningful, we must assume $\delta$ is small enough ($\delta\leq\sqrt{3/B_{w}B_{q}^{\prime\prime\prime}}$), so that:
@@
We note that we should consider two $g^{\star}$s, i.e., $q_{\beta}$ for $\beta\in\{\tau\pm\delta\}$. For notational convenience, let $(C_{\mathrm{norm}},\nu)$ be the constants in Assumption 5.1.3 When Assumption 5.1.3 is verified via the Hölder-type norm interpolation in Lemma B.6, one may take $C_{\mathrm{norm}}=C_{H}R^{d/(2s+d)}$ and $\nu_{H}=2s/(2s+d)$. To avoid confusion, we keep $\nu$ for the abstract exponent in Assumption 5.1.3 and use $\nu_{H}$ only for this Hölder instantiation.
@@
For classic statistical regression models, e.g., parametric model or Reproducing Kernel Hilbert Space (RKHS) regression with smooth kernel (e.g., Gaussian kernel), we have $\nu\approx 1$.
@@
$\hat{\Delta}_{\delta}(x)$ $\geq\Delta_{\delta}(x)-|\Delta_{\delta}(x)-\hat{\Delta}_{\delta}(x)|$ (68)
$\geq\Delta_{\delta}(x)-\frac{1}{2\delta}(|\hat{q}_{\tau+\delta}(x)-q_{\tau+\delta}(x)|+|\hat{q}_{\tau-\delta}(x)-q_{\tau-\delta}(x)|)$
$\geq\frac{1}{2B_{w}}-\frac{1}{2\delta}(\|\hat{q}_{\tau+\delta}-q_{\tau+\delta}\|_{L_{\infty}}+\|\hat{q}_{\tau-\delta}-q_{\tau-\delta}\|_{L_{\infty}})$
@@
$C_{\text{norm}}\left(C_{\text{fast}}\mathfrak{R}_{n}(\mathcal{G})\right)^{\nu}\leq\frac{\delta}{4B_{w}}.$ (69)
In fact, our optimal rate is given by $\delta^{\star}\asymp\mathcal{R}_{n}^{1/3}$, and thus we need $\nu\geq 1/3$, otherwise this stability condition would induce a slower rate. In the Hölder instantiation of Lemma B.6 where $\nu=\nu_{H}=2s/(2s+d)$, the condition $\nu\geq 1/3$ is equivalent to $d\leq 4s$, reflecting the standard curse of dimensionality. We remark that this bottleneck enlightens our clipping mechanism on the weights.
@@
Hence, we can applying Bernstein’s inequality (Bach, 2024) to transform the empirical norm into population norm, with probability at least $1-\zeta$:
@@ B.3.5 Optimal bandwidth and final bound
$c_{2}=4B_{w}^{2}\sqrt{4\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}}.$ (76)
@@
$U(\delta^{\star})/M_{\rho,\mathcal{G}}$ $=3\left(\frac{c_{1}}{4}\right)^{1/3}c_{2}^{2/3}$ (77)
$=C_{\mathrm{const}}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}\left(4\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}\right)^{1/3},$
@@
$+2M_{\rho,\mathcal{G}}C_{\mathrm{const}}B_{w}^{2}(B_{q}^{\prime\prime\prime})^{1/3}\left(4\mathcal{E}_{q}(n)^{2}+C_{\mathrm{norm}}\mathcal{E}_{q}(n)^{1+\nu}\sqrt{\frac{2^{3+2\nu}\log(1/\zeta)}{n}}+\frac{2C_{\mathrm{norm}}^{2}\mathcal{E}_{q}(n)^{2\nu}\log(1/\zeta)}{3n}\right)^{1/3}$
@@ C.1 Gap between $F_{S\mid X}(\hat{q}(x))$ and $\pi(x)$
In this part, we supplement the results proved in Plassier et al. (2025a).
@@
Here, we adopt a slightly stronger  [^3] We assume the existence of PDF, while in  (Plassier et al., 2025a), the continuity of CDF is assumed. version than (Plassier et al., 2025a) and suppress complex notations to enhance the readability. Assumption C.1 essentially characterizes the quality of quantile estimator, i.e., to what extent the quantile estimator removes the covariate-dependency on scores.
Theorem C.2 (Theorem 2 in Plassier et al. (2025a)).
@@
We note that the conformalization step is still substantial when $n$ is small (e.g., smaller than 100), while the $\Lambda$ would also decrease with the increase of the quality of the quantile estimator (and originally, the size of samples used for training the quantile estimator). Hence, with a reasonably large sample size of $\mathcal{D}_{\text{cal}}$, we can focus on characterizing the finite-sample performance of our method through analyzing the MSE of $F_{S\mid X}\left(\hat{q}_{\tau}(x)\right)$, and naturally the surrogate, expected risk of density-weighted pinball loss.